A rectangle has a perimeter of 44 cm. If the length is 6 cm longer than the width, find the dimensions of the rectangle.

Answer by checkley79(3341)

PERIMETER=2L+2W
L=W+6
44=2(W+6)+2W
44=2W+12+2W
44-12=4W
32=4W
32/4=W
8=W ANS.
L=8+6=14 ANS.
PROOF:
44=2*14+2*8
44=28+16
44=44